# Teleoscope: Exploring Themes in Large Document Sets By Example

### v1

@@ Abstract.
Qualitative thematic exploration of data by hand does not scale and researchers create and update a personalized point of view as they explore data. As a result, machine learning (ML) approaches that might help with exploration are challenging to apply. We developed Teleoscope, a web-based system that supports interactive exploration of large corpora (100K-1M) of short documents (1-3 paragraphs). Teleoscope provides visual programming workflows that have semantic and computational meaning; helping researchers to retrace, share, and recompute their sense-making process. Attempting to create qualitative “themes” rather than “topics,” our NLP approach tunes an ML model to “think like you” without significant retraining. Here, we present our two-year design process and validation of Teleoscope, including a multi-week study with qualitative researchers ($N=5$), a six-month field deployment with a qualitative research group, and an on-going public release.
datasets, machine learning, visualization
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXccs: Human-centered computing User interface programmingccs: Human-centered computing Field studies
@@ 1. Introduction
Exploring data at scale within a large corpus of documents is difficult, particularly if you want to interpret your data by telling a story or explaining a phenomenon. Qualitative research often focuses on interpretation, which we can think of as enriching data with context and meaning (Hennink et al., 2017; Sebele-Mpofu, 2020; LaDonna et al., 2021). For large corpora, summative statistical approaches are often used at the cost of interpretation. Qualitative research often focuses on interpretation by hand, which involves extensive reading to look for themes within data and connect empirical observations to theory or outside knowledge. But, working by hand is slow and limits the scale of what can be analyzed.
Machine learning (ML) is a statistical approach to data analysis; with the advent of large language models (LLMs), modeling context and meaning is becoming possible. A common summative ML approach is topic modelling: take an entire dataset and try to extract large-scale topics that label and partition document sets within a corpus. In contrast to topics, qualitative researchers aim to produce cross-cutting themes by enhancing existing, but perhaps non-obvious, connections between documents with explanation and theory.
LLMs use document similarity metrics for common semantic understandings (Devlin et al., 2018; Brown et al., 2020; Touvron et al., 2023; Cha, [n. d.]; OpenAI, 2023). To help qualitative researchers, a trained LLM would need to have the same understanding of document similarity as the researcher. However, training a LLM to think like one particular researcher would be impractical. Instead, in our work, we develop techniques and tools to refine LLMs through iteratively defining similarity metrics according to user interactions that are already part of the qualitative research workflow.
Good qualitative interpretation requires interaction with data to make sense of it. The “way we think” when making sense of data can be called a schema. The concept of a schema comes from psychology and refers to a cognitive framework that helps to organize and interpret particular events, behaviour and information. We might have many schemas that serve as frameworks to help interpret different events. These frameworks are difficult to introspect (i.e., know entirely through internal thoughts), but underlie our reactions to events in the world. For our purposes in this paper, the intuition, curiosity, and ad hoc hypotheses that researchers have about their data can be understood as expressions of their underlying schemas. An important part of the qualitative research process can be thought of as identifying, expressing and externalizing the researcher’s schema through labeling, arranging, modifying, and enriching data (Berret and Munzner, 2022). We can think of a schema as the internal cognitive structures and processes that produce our intuition and understanding that helps make sense of data. During research, as a person interacts with their dataset, they update and re-externalize portions of their implicit schemas and therefore updating their understanding. By externalizing, they are updating their explicit schemas. These processes feed back on each other.
To support thematic exploration of data (Li et al., 2023), it would be ideal if we could train an ML model that matches the user’s schema. But, training such an ML model is impractical with existing approaches: the schema is hidden, unique, and mutates every time the user interacts with data. Training and re-training such an ML model would therefore take a long time and require the user to generate many examples. Further, qualitative researchers often do not work alone; it is necessary to inspect, retrace, and share your thinking processes with other people on a research team.
To target this space between human and machine sense-making, we built Teleoscope, a system to support multiple users interactively exploring a large corpus based on LLM’s document similarity. With our system, users can iterate, collaborate, and group documents based on their own understandings of document similarities. Our system will then produce ‘fast enough’ and ‘good enough’ models by dimensionally reducing the LLM based on user-defined groups so that users can inspect and update their cognitive schemas alongside the ML models.
Problem: Thematic exploration of data by hand does not scale. But ML models cannot be easily used because they are not specific enough to capture a particular researcher’s schemas. Further, researchers need to update their schemas as they explore data, especially as they share findings with others.
We present Teleoscope, an interactive web-based system for exploring large datasets (100K-1M) of short documents such as social media posts. Teleoscope was designed to assist qualitative researchers during the discovery, data collection, and organization phases of a research study. Using Teleoscope, researchers can quickly search for themes across a document corpus, organize documents, and test ad hoc hypotheses before committing to a deeper analysis. Teleoscope’s interface is a visual data workflow editor that controls a cloud-based backend of machine learning (ML) workflows which calculate document similarity and clustering based on a user’s document grouping and workflow organization on the interactive workspace (see Fig. 1).
Figure 1. (1) Users start by performing a keyword search to explore documents; (2) Documents are dragged onto the workspace; (3) Documents can be put into groups for organization; (4) Teleoscope nodes can use documents, notes or groups as control inputs; (5) Projections create clusters using groups as control input; (6) Notes can contain arbitrary text which is also vectorized and can be used as a control input to a Teleoscope; (7) the sidebar has a quick viewer for documents, saved items, bookmarks, and settings. Keyboard navigation is included for quick exploration and group creation.
A key difficulty in using ML to support exploratory qualitative research is to define the right document similarity metric. This is because document similarity from a machine perspective can be quite different from a user’s understanding of document similarity (see Fig. 4). Teleoscopes’s natural language processing (NLP) system is a special case of semi-supervised topic modelling based on document similarity. However, the machine-generated outputs of topic modelling are often disappointingly general and miss the nuance that a researcher would prefer to have about their document groups (Nikolenko et al., 2017; Rüdiger et al., 2022). A researcher may decide that two documents that seem very different to the machine are part of the same theme. Or a machine may decide that two documents are very similar that a researcher would not think belong together. Teleoscope uses researcher’s understanding of similarity as expressed through grouping and mixing documents to define and express a distance metric that is used for dimensionality reduction and clustering.
Teleoscope is best thought of as a tool for thematic exploration rather than topic modelling. Our key insight is to continually update document representations based on user interaction that are good enough and fast enough to support iterative sense-making. We use the natural interactions of a qualitative researcher when interpreting documents such as arranging, annotating, and grouping. Teleoscope translates these interactions into inputs for a custom distance metric for ranking, dimensionality reduction and clustering. Put another way, the user starts from a large model and proceeds to refine it by (1) exploring documents in terms of keyword search; (2) discovering document similarities; (3) grouping documents together to indicate their personal understanding of similarity; (4) “mixing” them by connecting them to a teleoscope [^1] Our on-screen operation was unfortunately named for our tool early in the design process and has not been revised yet. In this paper, we use lower-case and italics to refer to the operation, and capitalize the system. Our apologies to the reader. operation to discover what the machine “thinks” is similar; (5) building up a workflow graph (used as input for dimensionality reduction); (6) iterating and refining input (see Fig 1).
We want to use the power of ML on large corpora, but human cognitive schemas are typically implicit. ML systems need data to be explicit, but making too many definitive choices too early can cut off important interpretive directions for a qualitative researcher (Chen et al., 2018). Iteratively making schemas explicit allows us to slowly approximate and develop a researcher’s schema by customizing an ML model to their research context.
We designed Teleoscope to make cognitive schema expression as explicit as possible by creating computable visual traces so that researchers could retrace, inspect, and modify their schemas at the same time as refining machine data representations. Effectively, we wanted to make a computable mind map that corresponds directly to an ML model state (an ML schema to match a cognitive schema). The word “Teleoscope” is a portmanteau of telescope and the Greek word telos (as in, teleology), which refers to explaining the purpose of something. Teleoscope is a way for researchers to intuitively explore and explain meaning and purpose for their datasets.
As such, our design goals targeted the creation of a system that allows continuous live interaction, as opposed to many topic modelling packages that run as python programs in the terminal (Pedregosa et al., 2011; Rehurek and Sojka, 2011; Devlin et al., 2019; El-Assady et al., 2019). Similarly, we wanted to a visualization system that captured the process of developing a thematic understanding of the data so that research collaborators could share their thought processes with each other. This is as opposed to topic modelling visualization systems that focus on high-level statistics, topic hierarchies, word counts, and labels (El-Assady et al., 2019; Kim et al., 2019; Terragni et al., 2021). Finally, we believe that the embodied experience of arranging documents is an important part of creating the sense of understanding. To support our target users—who are not computer experts—we wanted to make a web platform that allowed for fast enough interaction that users could feel as if they were having an embodied, creative, improvisational experience with the data. Our design goals are summarized here:
  • DG1. Create an interactive system where researchers directly manipulate data to produce ML models which then feed back into their understanding of the data. Imbue common qualitative interaction metaphors with computation; specifically those that correspond to actions that qualitative researchers are familiar with from data interpretation and exploration, such as iterative arranging, annotating, and grouping of documents. Qualitative researchers often use emergent interpretive processes such as physically re-arranging and annotating documents until patterns and meanings emerge.
  • DG2. Help researchers externalize their thought processes (cognitive schemas) as visual traces that can be retraced, modified, and inspected by collaborators. When interpreting data, capturing the steps of a thought process is useful when inspecting and changing your conclusions, and then later sharing them with others. If the visual representation traces your thought process but also has computational meaning, then changing the representation gives feedback to your schemas and vice-versa. Then, you can modify the representations when collaborating with other team members or reviewers to inspect, share and modify workflows and understandings.
  • DG3. Facilitate an on-going feeling of improvisational, creative, in-the-moment data exploration for researchers. The feeling of continual interactivity was important to us to facilitate through our system design choices. We wanted users to be able to continually interact with the system while inevitably large computations completed. We also wanted to limit the scope of computations to produce results that are good enough to inspect, but not perfect since we expect them to be quickly iterated on.
@@
To support our design goals and evaluate our system, we ran three field deployments. In the first deployment (N=5), we asked qualitative researchers to use Teleoscope to explore a shared topic over the course of a few weeks, then discuss their experience and findings at a focus group. For the second deployment, we recruited a three-person qualitative research team to use Teleoscope as part of their data collection and analysis process and responded to their design requests. For the third deployment, we released a live public version of Teleoscope which is currently running and available. This helped us to gain insight into our research goals.
At a high level, we were interested in seeing whether we had created a tool that would be effective in supporting qualitative researchers in the data collection phase, particularly in externalizing their cognitive schemas. We wanted to iteratively develop interaction and visual metaphors that capture these schemas and ensure that they are understandable to non-expert users. Once we developed the tool to a deployable level, we wanted to study how Teleoscope was used in a real collaborative research study.
@@
Although our main design inspiration and target demographic are qualitative researchers, our studies demonstrate that Teleoscope supports people in general-purpose text data exploration and discovery. We performed a software design process over two and a half years which was heavily informed by our own use of Teleoscope for qualitative analysis.
Our design goals, design studies, and software design process have produced the following contributions to the CHI community:
  • C1. Collaborative, interactive cognitive schema data exploration and sense-making. We contribute a system that concretely demonstrates a method for expressing and updating cognitive schemas alongside ML models as a form of collaborative data exploration in large corpora. We also provide our simple-yet-hard-to-develop interaction metaphors, assembly of ML approaches, and systems engineering approaches that facilitate the real-time interactive feeling of continuous improvisation required to make cognitive schema development possible alongside ML model interaction. To our knowledge, no other system has attempted to demonstrate a design for aligning machine and cognitive schemas, particularly at scale with a real-time system.
  • C2. Design studies of cognitive schema data exploration. We further contribute studies of our novel approach that detail both development and long-term field deployments of the system. Our studies are instructive for the community in that they detail a long a difficult design process for a simple outcome, and culminate in an unique long-term field deployment that involved a full real in-situ research use case. They further shed light on how qualitative methodologies can be understood with collaborative ML systems.
@@
  • C3. An open-source, usable, and adaptable product that can be collaboratively used by multiple team members. Teleoscope is provided as a research product that is both publicly deployed at [anonymized for review after consultation with CHI chairs] and is an open-source software for others to adapt and deploy themselves. Multiple users are able to collaborate in exploring over 500K documents on the same workspace; the deployment features user authorization, is cloud-native, works in the browser, and is robust enough to run continuously.
In this paper, we will address Related Work in Section 2 where we will explain the theoretical motivation for Teleoscope for qualitative methodologies, differentiate Teleoscope from other visualization and topic modelling tools, and explain the NLP approaches we are drawing from. We then explain the design of our system and motivate it with respect to qualitative researcher needs in Teleoscope Design in Section 3. After, we outline our participant-focused Design Studies in Section 4, wherein we performed an iterative low-cost design process, and three long-term field deployments to test and refine the viability of our approach. We discuss the results of our studies in light of our design goals and research questions in Section 5, and finish with a short Conclusion in Section 6.
2. Related Work
In this section, we outline the theoretical positioning of Teleoscope within the literature. We explain the relevant technologies, outline the theoretical connection to qualitative research, and motivate our design decisions relative to common approaches for creating interfaces for exploring large document sets and topic modelling. Teleoscope draws from three categories of related work: data visualization, the NLP field of topic modeling, and computer-supported sense-making as articulated through qualitative methodological theory and approaches.
As an interactive system, Teleoscope primarily takes inspiration from topic modelling visualization and qualitative analysis software. Similar to common topic modelling visualizations addressed below, Teleoscope structures and annotates a vast document space with labels and groups, supports creative and interactive exploration of documents along with document structures, and uses human-in-the-loop machine clustering of documents. We differentiate ourselves by taking a provenance-based approach in that we make the questions of “what did we find?” and “how did we get here?” the primary focus of manipulating objects in the interface (DG1). As a result, we further differentiate ourselves in terms of facilitating (1) collaborative sense-making (DG2); (2) at a large scale while remaining improvisational (DG3); and (3) targeting themes rather than topics.
@@ 2.1. What qualitative researchers need
Teleoscope’s design is largely inspired by the qualitative practice of thematic analysis (Braun and Clarke, 2012) and the methodological considerations therewith. This is well captured in the following quote by Nowell et. al:
@@
Teleoscope takes the same value propositions from the analysis phase and applies it to the data collection phase. Due to the human-in-the-loop ML modelling approach and the large amount of data, the collection methodology must also satisfy the standards of being precise, consistent, and exhaustive. However, at scale, there may be hundreds or thousands of documents that are thematically identical for a particular study. Researchers need to determine which documents are most similar within a corpus, structure the documents into groups, then find exemplar documents to construct a sample for more in-depth analysis. Teleoscope supports this data curation process by providing potential theoretical rigour that would apply standard thematic analysis by hand by focusing on visualizing the provenance of the ML model state (see below for a discussion on provenance).
@@
How then do we decide which documents are worth looking at? The Teleoscope approach prioritizes the sense-making process which is best understood by the concept of making tacit schemas explicit, as articulated by Berret and Munzner (Berret and Munzner, 2022). In quantitative methodologies, the standard approach is that scientists formulate their research as a series of falsifiable claims which are supported or rejected due to empirical evidence. In empirical qualitative methodologies, a researcher’s job is to interpret empirical evidence so that it makes sense for both the researcher and the reader. Berret and Munzner explain the act of sense-making as moving between tacit schemas (i.e., gut feelings, unconscious ideas) and explicit schemas (i.e., drawings, writings, and other data representations). By turning a tacit schema into one or more explicit schemas a researcher inspects their tacit schemas and makes them available for critique by others.
@@
Importantly, sense-making is an iterative process that starts from ambiguity. Chen et al. (Chen et al., 2018) argue that ML support for qualitative analysis in social science research needs to “[shift] the focus to ambiguity.” To paraphrase, computer scientists often create supposedly definitive and accurate ML models of textual data too early: before there has been a period of ambiguity. This is antithetical to qualitative definitions of academic rigour where it is considered important to preserve ambiguity for as long as possible. This allows for more time to make connections between data points and enrich theoretical insight while developing themes about the data. Good themes are rich (more interconnected) rather than accurate (categorically definitive) (Moser and Korstjens, 2017).
To balance the tension between making schemas explicit and retaining ambiguity, we designed Teleoscope to show users results that were “good enough” and “fast enough.” This motivated our choice to make an asynchronous processing system: users can continue to explore and arrange documents while waiting for a longer background ML calculation to complete. Since the act of arranging documents takes place within a node-based visual workflow editor, the arranging and linking of elements also has explicit computational meaning which is continually updated as the user “updates” their implicit cognitive schema.
@@
Teleoscope contributes to qualitative methodologies by importing standards of analysis into standards of document collection at scale (DG3). It also demonstrates a method for making the process of tacit sense-making explicit by asking the user to primarily manipulate workflows which represent their sense-making process (DG2). Unlike other systems, Teleoscope makes a schema expression an explicit goal, and provides a direct correspondence between the schema as expressed through visual workflow and ML model (DG1).
2.2. Visualization Approach
2.2.1. Provenance

### v2

# Crystallizing Schemas with Teleoscope: Thematic Curation of Large Text Corpora

@@ Abstract.
Making sense of large text corpora is difficult when scales reach thousands or millions of documents. With the advent of LLMs, the potential for large-scale sense-making is being realized. However, this presents a need for rigour in the data curation stage of thematic analysis: selecting the right documents to achieve appropriate information power (saturation) requires an auditable trace of researchers’ thought processes.
In this paper, we present methodological and design findings from a three-year design process where we worked with qualitative researchers to create an open-source platform called Teleoscope designed to rigorously curate documents at scale. By implementing the qualitative research values common to thematic analysis during the curation stage (which we call thematic curation), we found researchers could come to a shared understanding of a large corpus and feel confident in their curation decisions (which we call schema crystallization).
Qualitative research, Thematic curation, Schema nucleation, Schema externalization
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Collaborative and social computing systems and toolsccs: Human-centered computing Visualization systems and tools
Figure 1. A screenshot of the core Teleoscope workflow: starting from a keyword search, you choose documents to iteratively mix together, and then explore the ranked output documents.
@@ 1. Introduction
Exploring data at scale within a large corpus of documents is difficult, particularly if you want to interpret your data by telling a story or explaining a phenomenon. Qualitative research often focuses on interpretation, which we can think of as enriching data with context and meaning (Hennink et al., 2017; Sebele-Mpofu, 2020; LaDonna et al., 2021). For large corpora, quantitative statistical approaches are often used at the cost of qualitative interpretation. Interpretation is often done by hand, which involves extensive reading and annotating data while developing themes during a thematic analysis (Braun and Clarke, 2012). But, working by hand is slow and limits the scale of what can be analyzed.
For large document corpora (100K to 1M), some machine learning (ML) support is necessary due to the large volume of data. With the advent of large language models (LLMs), modeling context and meaning is becoming possible, providing the potential to automate interpretation.
However, a rigorous qualitative researcher cannot simply input data into ChatGPT and ask for a summary. Instead, researchers need to demonstrate the validity of their interpretations by showing the provenance of their analysis—to themselves, to their colleagues, and to reviewers. When analyzing, researchers aim to produce cross-cutting themes by drawing out related meanings between documents using interpretation and theory.
In the context of LLMs, “relatedness” is encoded as distance between vectors that represent documents (a.k.a., document embeddings) (Devlin et al., 2018; Brown et al., 2020; Touvron et al., 2023; Cha, [n. d.]; OpenAI, 2023). To help create themes, a LLM would need to represent vector similarity in a way that corresponds to how a researcher thinks documents are related. From a computational perspective, this means finding a way to represent underlying meanings and subtext using document embeddings.
Creating vector representations of meaning and subtext is no small task. Particularly while a researcher is actively trying to make sense of their data. In this paper, we are calling the “way we think” when making sense of data a schema (Berret and Munzner, 2022a), and support the thematic curation of data (Li et al., 2023) by creating a system that helps to track how a researcher thinks about their data, i.e., models their schema. However, schemas are hidden, unique, and change every time the user interacts with data. Further, qualitative researchers often do not work alone; it is necessary to inspect, retrace, and share your thinking processes with other people on a research team.
To enable collaborative human and machine sense-making, we built Teleoscope, a web-based system to support multiple users interactively exploring a large corpus of documents such as social media posts (100K-1M) based on LLM document similarity. Teleoscope assists qualitative researchers during the discovery, data curation, and organization phases of a research study. To understand one’s own thinking process, for purposes of rigor and sanity, researchers need to continually justify “how they got here,” not just to colleagues, but to themselves. We therefore focused our design efforts on visualizing data provenance to support rigorous and opinionated refining of a dataset, which we call thematic curation.
@@
To design and evaluate our system, we ran three field deployments. In the first deployment (N=5), we asked qualitative researchers to use Teleoscope to explore a shared topic over the course of a few weeks, then discuss their experience and findings at a focus group. For the second deployment, we recruited a three-person qualitative research team of nurses to use Teleoscope as part of their data curation and analysis process and responded to their design requests. For the third deployment, we released a live public version of Teleoscope which is currently running, available, and has active users who provide feature requests.
@@
Our studies demonstrate that Teleoscope supports professional researchers in general-purpose text data exploration and discovery. We performed a software design process over two and a half years which was heavily informed by our own use of Teleoscope for qualitative analysis. To summarize, our contributions are:
  • $\star$
    C1. Methodological analysis of thematic curation and a process we call schema crystallization. Teleoscope enables an extension of thematic analysis into document curation to enable more rigorous data interpretation where ideas about datasets are externalized, shared, and iterated upon.
  • $\star$
    C2. Results of case studies with qualitative researchers. Our long-term field deployments of the system demonstrate that qualitative methodologies can be supported with our approach of schema crystallization.
  • $\star$
@@
    C3. An open-source, usable, and adaptable product that supports collaboration. Teleoscope is provided as a research product that is publicly deployed (Teleoscope.ca, 2024a)and is open-source software (Teleoscope.ca, 2024b). Multiple users collaborate in exploring over 500K documents on the same workspace. The system is cloud-native, works in the browser, and is robust enough to run continuously.
2. Design Goals, Background, and Teleoscope’s Design
In this section, we outline our design goals and the relationship to existing literature. We explain relevant technologies, outline qualitative research theory, and motivate design decisions. We situate Teleoscope as a system for externalizing researcher’s internal schemas about their datasets.
The concept of a schema comes from psychology and refers to a cognitive framework that helps to organize and interpret particular events, behaviour, and information. For this paper, the intuition, curiosity, and ad hoc hypotheses that researchers have about their data can be understood as expressions of their schemas. An important part of the qualitative research process can be thought of as identifying, expressing and externalizing the researcher’s schema through labeling, arranging, modifying, and enriching data (Berret and Munzner, 2022a).
The word “Teleoscope” is a portmanteau of telescope and the Greek word teleology, which is about explaining the meaning and purpose of something. Teleoscope is a way for researchers to intuitively explore and explain meaning and purpose for their datasets. We use the natural interactions of a qualitative researcher when interpreting documents such as arranging, annotating, and grouping.
Our design goals target continuous live interaction, as opposed to topic modelling packages that run as terminal-based Python programs (Pedregosa et al., 2011; Rehurek and Sojka, 2011; Devlin et al., 2019; El-Assady et al., 2019). We visualize the process of thematic data exploration so that collaborators can share their thought processes and conclusions. This is different than topic modelling visualization systems that focus on high-level statistics, topic hierarchies, word counts, and labels (El-Assady et al., 2019; Kim et al., 2019; Terragni et al., 2021). To support our target users—who are not computer experts—we created a web platform to facilitate fast enough interaction that users could feel as if they were having a creative and improvisational experience with the data.
Our research design goals are:
  • DG1. Imbue common qualitative research interactions with computation. Qualitative researchers are familiar with arranging, annotating, and grouping of documents and use emergent interpretive processes such as physically re-arranging and annotating documents until patterns and meanings emerge.
  • DG2. Help researchers externalize their schemas as visual traces that can be retraced, modified, and inspected by collaborators. When interpreting data, capturing the steps of a thought process is useful in collaborating with team members to inspect, share and modify workflows and understandings.
  • DG3. Facilitate an on-going feeling of improvisational, creative, live data exploration for researchers. The feeling of continual interactivity was important to us to facilitate through our low-level system design choices, not having users be creatively blocked by waiting for long computations.
@@ 2.1. What qualitative researchers need
Teleoscope’s design is largely inspired by the qualitative practice and methodological values of thematic analysis (Braun and Clarke, 2012). This is well captured in the following quote by Nowell et. al:
@@
Teleoscope takes the same value propositions as a thematic analysis phase and applies it to the data curation phase, i.e., collected with an interpretive perspective (Mannheimer, 2021). Data curation must satisfy the thematic analysis standards of being rich, precise, consistent, and exhaustive. Information power (an evolution of theoretical saturation) is used in qualitative analysis as an analogue to statistical power in quantitative analysis (Malterud et al., 2016; Guest et al., 2020). At scale, there may be hundreds or thousands of documents that are thematically “identical” and provide no extra information power relative to a research question. Researchers therefore need to determine which documents are most similar within a corpus, structure the documents into groups, then find exemplar documents for more in-depth analysis. Teleoscope supports data curation by providing potential theoretical rigour that would apply standard thematic analysis by hand by focusing on visualizing the provenance of the ML model.
@@
How then do we decide which documents are worth looking at? In quantitative methodologies, the standard approach is that scientists formulate their research as a series of falsifiable claims which are supported or rejected due to empirical evidence. In qualitative methodologies, a researcher’s job is to interpret evidence so that it makes sense for both the researcher and the reader. Berret and Munzner explain the process of sense-making as moving between tacit schemas (i.e., gut feelings, unconscious ideas) and explicit schemas (i.e., drawings, writings, and other data representations) (Berret and Munzner, 2022a). By turning a tacit schema into explicit schemas, a researcher makes their tacit schema available for inspection and critique.
@@
Importantly, sense-making is an iterative process that starts from ambiguity. Chen et al. (Chen et al., 2018b) argue that ML support for qualitative analysis in social science research needs to “[shift] the focus to ambiguity.” To paraphrase, computer scientists often create supposedly definitive and “accurate” ML models of textual data too early, before there has been a period of ambiguity. Creating models too early is antithetical to qualitative definitions of academic rigour where it is considered important to preserve ambiguity for as long as possible. Having more time to make connections between data points can enrich theoretical insight while the researcher develops themes about the data. Good themes are rich (more interconnected) rather than accurate (categorically definitive) (Moser and Korstjens, 2017).
With Teleoscope, the act of arranging documents takes place within a node-based visual workflow editor. Arranging and linking elements onscreen both makes a schema explicit and runs the ML workflow. The computational schema updates as the user “updates” their cognitive schema. Braun and Clarke emphasize that during the data familiarization phase, there needs to be active engagement with the data:
{quoting}
Note-making helps you start to read the data as data. Reading data as data means not simply absorbing the surface meaning of the words on the page, as you might read a novel or magazine, but reading the words actively, analytically, and critically, and starting to think about what the data mean. (Braun and Clarke, 2012)
We consider Teleoscope to be a method of capturing the process of engaging with data that creates a correspondence between the externalized annotations on the screen, computational meaning with an ML workflow, and internal meaning for a researcher.
@@
Teleoscope contributes to qualitative methodologies by importing standards of data analysis to data curation at scale (DG3). It also demonstrates a method for making the process of tacit sense-making explicit by asking the user to primarily manipulate workflows which represent their sense-making process (DG2). Unlike other systems, Teleoscope makes a schema expression an explicit goal, and provides a direct correspondence between the schema as expressed through visual workflow and ML model (DG1).
Figure 2. An image of the Teleoscope workspace. (1) Users start by performing a keyword search to explore documents; (2) Documents are dragged onto the workspace; (3) Documents can be put into groups for organization; (4) Rank nodes can use documents, notes or groups as control inputs; (5) Projections create clusters using groups as control input; (6) Notes can contain arbitrary text which is also vectorized and can be used as a control input to a Rank; (7) the sidebar has a quick viewer for documents, saved items, bookmarks, and settings. Keyboard navigation is used for quick exploration and group creation.
2.2. Teleoscope’s Visualization Approach